OpenStudy (anonymous):

which is the correct partial fraction for 1/z^2(z+1) A. (Ax+B)/z^2+C/(z+1) B. A/z+B/Z^2+C/(z+1)

OpenStudy (bradely):

B is correct

OpenStudy (anonymous):

why is A not correct it is z^2

zepdrix (zepdrix):

\[\Large\rm \frac{1}{z^2(z+1)}=\frac{Az+B}{z^2}+\frac{C}{z+1}\]Neither option is incorrect. You can think of the z^2 as a `repeated linear factor` in which case you would break it down like in option B. Or you can think of it as a `quadratic` in which case it would break down with a linear factor on top as in option A.

OpenStudy (anonymous):

ohh thanks .. i was also confused about this ... i felt both are correct too.

zepdrix (zepdrix):

Option B is a little nicer to work with I think. Maybe that's just preference though :)